# On the Robustness of Fully-Spiking Neural Networks in Open-World Scenarios using Forward-Only Learning Algorithms

### v1

@@ 5.2 RQ2: Out-of-Distribution Detection Results
The metrics of the OoD detection experiments using our proposed method are presented in Table 2. After thoroughly analyzed the results, we conclude that they have yielded a positive outcome to our initial research question. A detailed discussion of these results follows in the present subsection.
Table 2: Results of the FF-SCP algorithm of our three proposed networks: the Analog FF-Network; the Unbounded FF-Network, denoted as $G_{\infty}(\cdot)$; and the Bounded FF-Network, denoted as $G_{0}(\cdot)$. Data from the SCP and ODIN has been extracted from the paper [24]. The ANN results are only shown as a baseline to compare ANN vs SNN.
@@
ID Dataset OOD Dataset FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN FF-SCP Analog FF-SCP $G_{\infty}(\cdot)$ FF-SCP $G_{0}(\cdot)$ SCP ODIN
@@
sFF-SCP is competitive against SCP and ODIN
@@
The main hypothesis guiding the development of the FF-SCP algorithm was that the improved representation capabilities of networks trained with FFA, together with refined geometric methods, would yield higher accuracy in OoD Detection tasks. The results, as presented in Table 2, verify this hypothesis, with both the bounded and unbounded models consistently outperforming the SCP algorithm across all datasets and metrics, except for one instance. The results are particularly evident when compared against the unbounded versions of the FF-SCP, where the average AUROC is $7.95$ points higher. Furthermore, the proposed algorithm exhibits a significant reduction in the FPR95, implying that it is more robust in situations where a high True Positive Rate is required. Nevertheless, the FF-SCP algorithm appears to follow the same pattern of the best and worse-performing pairs of ID and OoD datasets as the SCP algorithm, confirming the similarity between the heuristics of both methods. Both algorithms demonstrate the capability to outperform ODIN in the MNIST, FMNIST, and Letters datasets. However, they lack accuracy when used against the KMNIST dataset, where ODIN manages to outperform in several instances. A plausible explanation for the reduced results on the KMNIST as an ID dataset may arise from the suboptimal test performance exhibited by networks using FFA. Upon examining the weights they generate, they demonstrate less specificity, allowing for easier OoD samples to exhibit similar latent vectors to ID samples. Additionally, consistent with the findings of Martinez-Seras et al., the Omniglot dataset remains the poorest performing as an OoD dataset, with ODIN outperforming both algorithms [24].
Bounded vs Unbounded Goodness in FF-SCP
@@
Both goodness functions show remarkable similarity across most tasks in the benchmark, with the highest accuracy alternating depending on the specific task. However, the bounded goodness displays suboptimal accuracy in the KMNIST task, with numerous outcomes falling below 70 in terms of AUROC. This performance discrepancy may be attributed to the fact that the latent vectors of the bounded network lack sparsity, resulting in less specialization within the representation. This lack of neural specialization, together with the non-competitive accuracy on the dataset, may be explained by class clusters not achieving enough distance between them, thereby explaining the low accuracy when using the FF-SCP algorithm.
SNNs compete with ANNs when employing FF-SCP
@@
To ensure that SNNs offer practical advantages over ANNs, it’s essential to verify that their performance is not significantly overshadowed by their analog counterparts. Such a scenario would suggest that the architectural benefits of SNNs, such as inference speed and energy efficiency, might become secondary to accuracy advantages. The results presented in Table 2 demonstrate how the FF-SCP employed with unbounded goodness on SNNs yields outcomes comparable to its analog counterpart. Both methods achieve AUROC accuracies with an average difference of less than one point. Analog networks using FF-SCP perform slightly better than SNNs in average AUROC and both average and median AUPR. Conversely, both SNN goodness functions achieve superior results in median AUROC and both FPR95 measures. Examining specific instances, the worst result in analog networks is approximately 5 points below the worst result in its unbounded spiking counterpart. The bounded SNN counterpart experiences an additional drop when presented with KMNIST as the ID dataset, primarily due to the previously mentioned lack of test accuracy. However, the overall difference in accuracy is minimal, with this bounded version achieving higher AUROC in multiple cases. Nevertheless, the discrepancies between both spiking algorithms are minimal, suggesting that they offer similar results.
@@
As detailed in Subsection 3.2, the FF-SCP algorithm requires an additional conditional branch to ensure that the ordering in the scoring function is not reversed when the condition specified in Equation (13) is satisfied. To empirically demonstrate the necessity of this modification, the results obtained on the same benchmarks using the FF-SCP algorithm with the scoring function $s(x)$ are presented in Table 3, in Appendix C. In these results, the impact of this modification becomes evident when the ID dataset is KMNIST and the OoD dataset is FMNIST, where the AUROC drops significantly to 32.95 under models trained with the unbounded sFFA and to 17.92 for analog networks. However, when the condition specified in Equation (13) does not hold, the algorithm remains capable of accurately detecting OoD samples. This is particularly observed in the bounded version of the goodness, where the latent space exhibits highly active latent vectors, consistently avoiding the reversal in the scoring values.
Statistical significance
The obtained accuracy results across distinct models exhibit high similarity, posing a challenge in establishing clear evidence of the statistical significance of their outcomes. To address this, we rely on two widely used statistical methods. First, a Critical Difference Diagram [40], which incorporates a post-hoc Nemenyi test to calculate the critical distance value. This diagram orders various techniques based on their average rank, helping visualize the differences. If the average rankings of two methods are separated by less than the critical difference value, they are considered statistically equal; otherwise, they are considered statistically different.
Secondly, Bayesian posterior analysis measures the paired performance scores obtained from two different techniques. In their work [41], Benavoli et al. use a Monte Carlo sampling technique to estimate the likelihood of one technique outperforming another. This analysis distinguishes three outcomes: two indicating algorithmic superiority and one reflecting practical equivalence when the methods are statistically indistinguishable. The practical equivalence is adjusted by a hyperparameter called rope, which establishes the minimum difference between technique scores to consider their performance statistically different.
The first result from the Critical Distance Diagram is illustrated in Figure 4. This diagram provides evidence of the statistical equivalence between both versions of FF-SCP and ODIN. Additionally, the analysis is consistent with the results from Table 2, showing that the representational advantages of FFA leveraged in the FF-SCP algorithm result in a statistical advantage over the SCP algorithm.
[FF-SCP vs SCP]Refer to caption    [FF-SCP vs SCP]Refer to caption [FF-SCP vs ODIN]Refer to caption
Figure 4: a) Critical Distance (CD) diagram computed over the AUROC results across the cases considered in the Tables 2 and 3. The value of CD is shown on top of the diagram. Horizontal lines connecting the average rank markers of several detectors indicate that their average rankings cannot be claimed to be different to each other with statistical significance. b) & c) Bayesian Posterior Signed Tests computed over the paired AUROC scores obtained from the FF-SCP algorithm on the Unbounded sFFA against (b) the SCP algorithm and (c) ODIN. The rope is set equal to AUROC $=1\%$. We denote by sFF-SCP to the application of the algorithm under the latent space of the SNN using the unbounded goodness.
The second statistical method, Bayesian Posterior Analysis, has been depicted in Figure 4 and Figure 4. The first diagram demonstrates consistent results with the previous method, indicating that the use of FFA yields superior latent representations, resulting in improved OoD capabilities. Similarly, the results in Figure 4 reveal that the FF-SCP algorithm generally outperforms ODIN, with only a small subset of cases showing better performance for ODIN. These cases align with the observation that ODIN was able to perform better in some experiments, particularly with Omniglot as the OoD dataset and KMNIST as the ID dataset. Both algorithms exhibit almost no points within the rope region, implying that none of the methods perform equivalently, with each algorithm demonstrating statistical significance compared to the other.
@@ 5.3 RQ3: Attribution Results
[MNIST-Square with $\alpha=0.1$.]Refer to caption          [MNIST-Gaussian with $\alpha=0.7$.]Refer to caption
[MNIST-StripeOff with $\alpha=11$.]Refer to caption          [MNIST-StripeOn with $\alpha=0.5$.]Refer to caption
@@
Figure 5: Batch of four visually modified samples from the MNIST dataset with their reconstructions and the corresponding attribution maps, featuring a distinct visual obstruction: (a) Square; (b) Gaussian; (c) StripeOff; (d) StripeOn.
@@
To address the third research question and verify the quality of the attribution maps generated by our approach, as detailed in Algorithm 2, we have employed a qualitative analysis of several visually modified samples from the MNIST and KMNIST datasets. Recalling from the approach detailed in Section 3.3, our method provides an interpretability map by highlighting the regions of the image that push the original sample into the OoD space. Cold regions, represented by blue, indicate areas in the original input sample where certain features are missing compared to ID data. Conversely, hot regions, denoted by red, highlight areas in the original image that are not present in the ID data. The selected samples from the modified MNIST and KMNIST datasets are presented in Figure 5 and Figure 6 respectively.
One initial result, consistent with the findings from [5], is that the latent space provides sufficient representation capabilities to offer accurate reconstructions without excessively losing the overall original shape. Similarly, the results of the attribution on the MNIST samples appear to provide a positive answer to RQ3, offering clear attribution maps that directly highlight the incorporated artifacts. For instance, the clearest example can be observed in Figure 5, where the included square is clearly depicted as an artifact with a high temperature, demonstrating that the primary factor pushing this sample into the OoD region is the artifact itself. A similar case is seen in Figure 5, where the stripe is successfully localized with a cold temperature surrounding the missing area of the digit 1, showcasing the algorithm’s ability to detect instances where the figure has been split and to create a connected reconstruction. Conversely, Figure 5 has a horizontal stripe embedded into the image, which the algorithm detects and erases, preserving the prior regions of the number that overlap with the stripe. The final selected example is Figure 5, where the algorithm demonstrates its ability to reduce noise while preserving the shape of the digit. However, a noticeable change in the upper part of the number is apparent, which may arise due to the uncommon trace compared to the most representative samples of the class, which tend to have a closed loop instead of the open end.
Among the selected examples, the attribution map that best highlights the artifact is present in Figure 5. This additional precision, compared to the other samples, can be attributed to the lack of overlap between the original digit and the artifacts, allowing the algorithm to solely highlight the altered regions. In contrast, all other samples present slight alterations along the border of the digits, which do not appear to affect the topology of the digit. However, the degree of this deformation is not significant enough for a human evaluator to misinterpret the artifacts with respect to the ID data, so the results remain easily interpretable. For instance, in its most extreme case, as observed in Figure 5 with the StripeOn modification, the artifact region remains the most noticeable area, easily distinguishable from the noisy contour of the number.
[KMNIST-Square with factor of $0.1$.]Refer to caption           [KMNIST-Gaussian with factor of $0.7$.]Refer to caption
[KMNIST-StripeOn with factor of $0.5$]Refer to caption          [KMNIST-StripeOff with factor of $11$]Refer to caption
Figure 6: Batch of four visually modified samples from the KMNIST dataset with their reconstructions and the corresponding attribution maps, featuring a distinct visual obstruction: (a) Square; (b) Gaussian; (c) StripeOff; (d) StripeOn.
@@
As pointed out in the training results (see Subsection 5.1) and further confirmed in the OoD Detection results (see Subsection 5.2), the principal weakness of FFA becomes apparent in the KMNIST dataset, where the presence of complex features from the classes hinders the network’s representation capabilities, impacting both accuracy and OoD detection results. As a consequence, a sharp drop in accuracy within the attribution maps can be observed. This becomes evident when comparing results between the MNIST dataset and the KMNIST dataset, where the shape of KMNIST reconstructions is less accurate than those obtained from the MNIST dataset, resulting in noisier attribution maps. Nevertheless, the overall shape of the figure is still consistently reconstructed in most of the samples that are correctly classified.
@@
One such instance is observed in Figure 6, where the upper and lower limits of the character have been distorted, while the stripe artifact is properly removed and highlighted. A similar effect is present in Figure 6, where our algorithm can properly detect the embedded square, showing the highest temperature, while the resulting character outline has been slightly modified but remains within the original class distribution. In contrast, a greater shift can be observed when the modifications to the image are more drastic. Such is the case for Figure 6, where the obtained reconstruction is able to fill the missing stripe, but the contour appears blurred and several previously sharp features are missing, such as the upper stroke. Finally, in Figure 6, the algorithm is able to remove a significant portion of the noise with minimal changes to the shape of the characters and only a small amount of blur. Although all samples exhibit slight perturbations and do not present clean attribution, the discriminative regions of the embedded artifacts can be clearly detected. Therefore, our results provide a positive answer to RQ3, showcasing the capability of our algorithm to create interpretable attribution maps that explain the features pushing the samples into the OoD regions. However, it is important to acknowledge the limitations that occur with suboptimal models in terms of accuracy, whose reduced representational capabilities prove detrimental to the subsequent OoD and interpretability techniques.
$\bm{\alpha}$ Value Exploration
The only hyperparameter of our method is the $\alpha$ value, which balances the trade-off between the distance to the class cluster and the distance to the initial image. Higher values of $\alpha$ bring the reconstruction closer to samples from the ID data, while lower values result in more similar reconstructions to the input images. This effect is exemplified in Figure 7, where we present several reconstructions of the sample from Figure 5 with different values of $\alpha$. In the different pictures, we can observe how lower values of this hyperparameter can generate overfitted reconstructions that are not sampled from the ID cluster, keeping the digit split in two. By increasing the value of this hyperparameter, it allows for less overfitted reconstructions, connecting the two components into the corresponding number.
Since the network trained on the MNIST dataset contains accurate representations of the values, in several instances, the alpha value does not need to heavily prioritize the distance to the cluster. This effect is clear in Figure 5, where a value of $0.1$ is enough to accurately highlight the artifact in the heatmap. Conversely, there are instances where higher alpha values are required for accurate reconstructions, notably in cases where the complexity of the initial image or the artifact is higher. An example can be observed in the KMNIST dataset in Figure 6, where, in order to properly remove the stripe from the image, higher values of alpha are required. This increase creates a sample with more distance from the original sample but increases the distance to the ID. In this case, the reconstruction is close enough to understand that the main artifact in the image is the horizontal stripe added to it.
[$\alpha=0.1$]Refer to caption        [$\alpha=1$]Refer to caption        [$\alpha=11$]Refer to caption
Figure 7: Reconstructions of the sample from Figure 5 with different values of $\alpha$.
6 Conclusions and Future Work
The field of Artificial Intelligence has experienced rapid advancements over the last decade, even surpassing human experts in multiple tasks. However, the development of methods to ensure the robustness of these models has not kept pace with those driving performance improvements. This imbalance has resulted in the accuracy of models being overshadowed by their black-box nature when integrated into production-ready software. This issue is particularly critical in neuromorphic computing, where the temporal dynamics of spikes pose a challenge in analyzing model behavior. Nonetheless, the introduction of the Forward-Forward Algorithm (FFA) has offered a promising biologically plausible alternative learning mechanism, offering a latent space with highly representative properties This spaces are characterized by a well-defined distribution of latent vectors that cluster with other samples from the same class while remaining distinct from instances of other classes.
This work proposes an adaptation of FFA for spiking neural networks by developing two alternative formulations of the goodness functions and utilizing surrogate gradients to circumvent the non-differentiability of the neural models. Capitalizing on the representational properties of these networks, we develop a novel Out-of-Distribution (OoD) detection algorithm, denoted as FF-SCP, which uses the distances between latent vectors and the latent manifolds of classes as an effective measure of the likelihood of a sample belonging to the training distribution. To enhance the functionality of this method, we introduce a novel attribution method that explores the latent space to generate an attribution map, highlighting features in the sample that differ from an arbitrary class. To verify the competitiveness of our FFA adaptation, we trained the network on multiple well-known datasets (e.g., KMNIST, EMNIST) and used these networks to evaluate the performance of our FF-SCP algorithm when presented with OoD samples from other conventional baselines (e.g., Omniglot, NotMNIST). Additionally, we performed a qualitative analysis on our interpretability algorithm by manually including artifacts into samples from the training data and then evaluating the quality of the generated attribution maps. Our results provide additional evidence of FFA’s capability to deliver competitive results in spiking neural networks compared to its analog formulation. Furthermore, we demonstrate how FFA can effectively detect OoD instances in spiking neural networks, outperforming SCP and ODIN. Similarly, a qualitative analysis on the distorted images shows how our method can accurately highlight the regions of the artifacts, allowing human evaluators to identify the regions pushing samples outside the image distribution. With these two algorithms, FFA proves to offer highly functional properties, resulting in a more explainable model than backpropagation.
The positive outcomes of our work pave the way for multiple future lines of research into the applicability of FFA. One clear line of research involves extending the key concepts from this work into the study of concept drift, which is critical in highly dynamic scenarios. Similarly, leveraging the findings on OoD detection by incorporating incremental learning techniques is crucial for implementing an Open-World Learning algorithm. It is also important to note that our results are limited by the lack of alternative architectures, such as convolutions or transformers, which could vastly improve accuracy in complex data domains. Consequently, one of our primary focuses will be adapting FFA to accommodate more complex dynamics, potentially enhancing performance in diverse data regimes.
Acknowledgments
This work has received funding support from the Basque Government through a number of grants: projects BEREZ-IA (KK-2023/00012) and CERBERO (KK-2024/00022) from the ELKARTEK program, a BIKAINTEK PhD grant (008-B2/2021) and the consolidated research groups Deusto for Knowledge - D4K (IT1528-22) and MATHMODE (IT1456-22).
Appendix A Reversal of Order in FF-SCP

### v2

@@ 5.2 RQ2: Out-of-Distribution Detection Results
The metrics of the OoD detection experiments using our proposed method are presented in Table 2 for dense models and in Table 3 for image data. After thoroughly analyzing the results, we conclude that they have yielded a positive outcome to our initial research question. However, we remark the current limitations of FFA when dealing with image data, which we analyze separately from the fully FFA-based implementation. A detailed discussion of these results follows in the present subsection.
Table 2: Results of the FF-OoD algorithm of our three proposed networks: the Analog FF-Network; the Unbounded FF-Network, denoted as $G_{\infty}(\cdot)$; and the Bounded FF-Network, denoted as $G_{0}(\cdot)$. Data from the SCP and ODIN has been extracted from the paper Martinez-Seras et al. (2023). The ANN results are only shown as a baseline to compare ANN vs SNN.
@@
ID Dataset OOD Dataset FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN
@@
sFF-OoD is competitive against SCP and ODIN
The main hypothesis guiding the development of the FF-OoD algorithm was that the improved representation capabilities of networks trained with FFA, together with refined geometric methods, would yield higher accuracy in OoD Detection tasks. The results, as presented in Table 2 and Table 3, verify this hypothesis, with both the bounded and unbounded models consistently outperforming prior methods in densely connected networks, and unbounded methods achieving higher AUROC’s in models with a convolutional backbone.
@@
Nevertheless, the FF-OoD algorithm appears to follow the same pattern of the best- and worst-performing pairs of ID and OoD datasets as the SCP algorithm, confirming the similarity between the heuristics of both methods. A plausible explanation for the reduced results on KMNIST as an ID dataset may arise from the suboptimal test performance exhibited by networks using FFA. Upon examining the weights they generate, they demonstrate less feature-selectivity, allowing for easier OoD samples to exhibit similar latent vectors to ID samples. Additionally, consistent with the findings of Martinez-Seras et al. (2023), the Omniglot dataset remains the poorest performing as an OoD dataset, with ODIN outperforming both algorithms. However, both FF-OoD and SCP demonstrate the capability to outperform ODIN in the MNIST, FMNIST, and Letters datasets, reinforcing their potential in specific scenarios.
Bounded vs Unbounded Goodness in FF-OoD
@@
Both goodness functions show remarkable similarity across most tasks in the benchmark, with the highest accuracy alternating depending on the specific task. However, the bounded goodness displays suboptimal accuracy in the KMNIST task, with numerous outcomes falling below 70 in terms of AUROC. This performance discrepancy may be attributed to the fact that the latent vectors of the bounded network lack sparsity, resulting in less specialization within the representation. This lack of neural specialization, together with the non-competitive accuracy on the dataset, may be explained by class clusters not achieving enough distance between them, thereby explaining the low accuracy when using the FF-OoD algorithm.
SNNs compete with ANNs when employing FF-OoD
@@
To ensure that SNNs offer practical advantages over ANNs, it is essential to verify that their performance is not significantly overshadowed by their analog counterparts. Such a scenario would suggest that the architectural benefits of SNNs, such as inference speed and energy efficiency, might become secondary to accuracy advantages. The results presented in Table 2 demonstrate how the FF-OoD employed with unbounded goodness on SNNs yields outcomes comparable to its analog counterpart. Both methods achieve AUROC accuracies with an average difference of less than one point. Analog networks using FF-OoD perform slightly better than SNNs in average AUROC and both average and median AUPR. Conversely, both SNN goodness functions achieve superior results in median AUROC and both FPR95 measures. Examining specific instances, the worst result in analog networks is approximately 5 points below the worst result in its unbounded spiking counterpart. The bounded SNN counterpart experiences an additional drop when presented with KMNIST as the ID dataset, primarily due to the previously mentioned lack of test accuracy. However, the overall difference in accuracy is minimal, with this bounded version achieving higher AUROC in multiple cases.
Table 3: Results of the FF-OoD algorithm of our three proposed networks: the Analog FF-Network; the Unbounded FF-Network, denoted as $G_{\infty}(\cdot)$; and the Bounded FF-Network, denoted as $G_{0}(\cdot)$. Data from the SCP and ODIN has been extracted from the paper Martinez-Seras et al. (2023). The ANN results are only shown as a baseline to compare ANN vs SNN.
AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$
ID Dataset OOD Dataset FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN FF-OoD Analog FF-OoD $G_{\infty}(\cdot)$ FF-OoD $G_{0}(\cdot)$ SCP ODIN
CIFAR10 Food101 85.25 84.02 76.61 71.65 68.10 91.39 91.27 86.17 70.30 70.12 48.57 59.02 66.46 81.56 89.99
Flowers102 86.88 85.64 79.44 75.77 68.71 78.08 79.95 63.66 75.04 70.93 48.57 60.61 52.63 75.59 90.54
Caltech101 70.63 68.48 67.79 67.02 65.88 65.94 63.76 60.46 64.47 67.90 83.44 81.30 81.39 83.81 91.53
GTSRB 89.76 88.82 81.27 78.76 71.81 89.55 89.99 79.72 78.10 75.42 28.08 50.69 52.21 69.49 91.38
SVHN 89.25 86.28 77.74 62.70 70.10 94.24 93.63 87.46 61.21 72.27 31.69 62.34 62.23 88.51 89.41
EuroSAT 90.27 87.98 75.10 65.88 70.00 95.37 95.14 86.74 64.91 71.43 33.98 62.41 63.49 85.69 88.96
DTD 83.74 82.48 77.13 61.43 81.76 45.80 53.86 33.43 63.91 85.12 62.50 71.69 66.10 90.60 83.95
SVHN Food101 81.48 85.59 70.89 61.73 63.97 71.83 77.82 64.46 59.24 64.94 45.73 36.38 76.56 84.89 92.71
Flowers102 83.66 87.24 72.86 80.47 77.92 43.47 53.49 32.78 81.66 82.68 36.64 33.81 68.88 60.79 93.20
Caltech101 79.19 83.54 69.22 80.94 77.24 45.93 53.60 36.89 82.75 81.86 57.67 47.18 77.83 60.35 91.97
GTSRB 80.72 81.04 56.30 81.82 77.02 60.55 64.83 36.41 82.23 81.85 56.61 67.03 87.21 56.24 90.09
CIFAR10 79.41 85.48 72.67 83.11 81.01 48.17 57.88 43.61 83.80 81.10 54.23 37.06 73.70 55.59 88.64
EuroSAT 82.49 84.18 73.49 79.06 77.73 78.32 83.03 68.19 79.38 82.25 53.10 60.00 63.13 60.84 92.32
DTD 77.47 82.28 70.60 81.16 77.69 13.56 17.30 11.83 92.85 82.39 55.90 43.73 73.83 56.75 81.61
Average 82.87 83.78 72.93 73.67 73.49 65.87 69.68 56.55 74.27 76.44 49.76 55.23 68.97 72.19 89.73
Median 83.07 84.83 73.17 77.26 74.41 68.88 71.32 62.06 76.57 78.26 50.83 59.51 67.67 72.54 90.31
@@
To empirically demonstrate the necessity of this modification, the results obtained on the same benchmarks using the FF-OoD algorithm with the scoring function $s(x)$ are presented in Table 4, in B. In these results, the impact of this modification becomes evident when the ID dataset is KMNIST and the OoD dataset is FMNIST, where the AUROC drops significantly to 32.95 under models trained with the unbounded spiking FFA and to 17.92 for analog networks. However, when the condition specified in Equation (13) (namely, the low intra-class distance) does not hold, the algorithm remains capable of accurately detecting OoD samples. This is particularly observed in the bounded version of the goodness, where the latent space exhibits highly active latent vectors, consistently avoiding the reversal in the scoring values.
FFA requires good image backbones
In contrast to the fully dense model presented in Table 2, the model employing a convolutional backbone shows mixed results. As shown in Table 3, its AUROC performance remains comparable to that of SCP and ODIN. However, the AUPR scores drop significantly when evaluated on the SVHN ID dataset. This discrepancy likely stems from the pretraining of ResNet18 on ImageNet, which shares greater similarity with CIFAR-10, another natural image dataset. In contrast, SVHN, which comprises images of numerical digits, differs significantly in distribution, leading to suboptimal AUPR performance compared to CIFAR-10. Nevertheless, the latent vectors contain enough feature information for FF-OoD to still high OoD Detection accuracy on instances where SVHN and OoD datasets are distinct enough (e.g. Food101 or EuroSAT).
@@ 5.3 RQ3: Attribution Results
(a) MNIST-Square ($\alpha=0.1$)
(b) MNIST-Gaussian ($\alpha=0.7$)
(c) MNIST-StripeOff ($\alpha=11$)
(d) MNIST-StripeOn ($\alpha=0.5$)
(e) KMNIST-Square ($\alpha=0.1$)
(f) KMNIST-Gaussian ($\alpha=0.7$)
(g) KMNIST-StripeOn ($\alpha=0.5$)
(h) KMNIST-StripeOff ($\alpha=11$)
@@
Figure 4: Batch of four visually modified samples from the MNIST and KMNIST datasets with their reconstructions and the corresponding attribution maps, featuring a distinct visual obstruction: a) Square; b) Gaussian; c ) StripeOff; d) StripeOn;e) Square; f) Gaussian; g) StripeOff; h) StripeOn.
@@
To address the third research question and verify the quality of the attribution maps generated by our approach, as detailed in Algorithm 2, we have employed a qualitative analysis of several visually modified samples from the MNIST and KMNIST datasets. Recalling from the approach detailed in Section 3.3, our method provides an interpretability map by highlighting the regions of the image that push the original sample into the OoD space. Cold regions, represented by blue, indicate areas in the original input sample where certain features are missing compared to ID data. Conversely, hot regions, denoted by red, highlight areas in the original image that are not present in the ID data. The selected samples from the modified MNIST and KMNIST datasets are presented in Figure 4.
One initial result, consistent with the findings from (Ororbia and Mali, 2023), is that the latent space provides sufficient representation capabilities to offer accurate reconstructions without excessively losing the overall original shape. For instance, in Figure 4a, the included square is clearly depicted as an artifact with the expected high temperature. Similarly, the horizontal stripe embedded in Figure 4d is properly detected and erased in its respective reconstruction. Analogously, Figure 4c showcases the algorithm’s ability to detect missing regions, exposing the missing stripe using cold temperatures on the attribution map. Finally, in a mixed example from Figure 4b, the algorithm detects the induced noise, providing a clean reconstruction of the sample. However, while the algorithm reports good quality attribution maps, all instances show slight distortions on the digit’s borders. Nevertheless, the degree of deformation does not appear to affect the topology of the digit and is not significant enough for a human evaluator to misinterpret the artifacts with respect to the ID data. For instance, in its most extreme case, as observed in Figure 4d with the StripeOn modification, the artifact region remains the most noticeable area, easily distinguishable from the noisy contour of the number.
@@
As pointed out in the training results and further confirmed in the OoD Detection results, the principal weakness of FFA becomes apparent in the KMNIST dataset, where the presence of complex features from the classes hinders the network’s representation capabilities, impacting both accuracy and OoD detection results. As a consequence, a sharp drop in accuracy within the attribution maps is observed. This becomes evident when comparing results between the MNIST dataset and the KMNIST dataset, where the shape of KMNIST reconstructions is less accurate than those obtained from the MNIST dataset, resulting in noisier attribution maps. Nevertheless, the overall shape of the figure is still consistently reconstructed in most of the samples that are correctly classified.
@@
One such instance is observed in Figure 4g, where the upper and lower limits of the character have been distorted, while the stripe artifact is properly removed and highlighted. A similar effect is present in Figure 4e, where our algorithm can properly detect the embedded square, showing the highest temperature, while the resulting character outline has been slightly modified but remains within the original class distribution. In contrast, a greater shift can be observed when the modifications to the image are more drastic. Such is the case for Figure 4h, where the obtained reconstruction is able to fill the missing stripe, but the contour appears blurred and several previously sharp features are missing, such as the upper stroke. Finally, in Figure 4f, the algorithm is able to remove a significant portion of the noise with minimal changes to the shape of the characters and only a small amount of blur. Although all samples exhibit slight perturbations and do not present clean attribution, the discriminative regions of the embedded artifacts can be clearly detected. Therefore, our results provide a positive answer to RQ3, showcasing the capability of our algorithm to create interpretable attribution maps that explain the features pushing the samples into the OoD regions.
5.4 Limitations observed in the experiments
FFA shows promise in enabling forward-only learning, particularly in datasets with high inter-class distance. This is evident in its strong performance on datasets such as MNIST, KMNIST, and EMNIST. However, when applied to image-based data using convolutional architectures, early experiments suggest that the algorithm struggles to converge to optimal solutions. It often either significantly underperforms or fixates on predicting a single class. This limitation makes applying FFA to natural image datasets like CIFAR-10 particularly challenging. In our work, we investigate the use of ResNet18 as a backbone, incorporating FFA as a secondary projection network to reshape the latent space into a more manageable geometry. Our results demonstrate that FFA can still effectively distinguish between ID and OoD data, even when the backbone is not explicitly trained for a specific task but instead derived from a pretrained general image dataset. However, the need for additional networks partially undermines the method’s independence, underscoring the necessity for further research into alternative implementations of FFA to enhance its viability within convolutional architectures.
6 Conclusions
This work introduces FF-OoD, a novel Out-of-Distribution (OoD) detection method for Spiking Neural Networks (SNNs) that employs the characteristic topology emerging from networks trained using the Forward-Forward Algorithm (FFA). To achieve this, we extend the original formulation of FFA’s goodness functions to accommodate the additional temporal dimension inherent to SNNs and study the novel cluster structure that arises among the so-called negative data. Given the significant separation observed between distinct classes within the latent space of these networks, FF-OoD computes the distance between input samples and precomputed representative manifolds of known classes. In addition, we provide theoretical insights into how high intra-class variability affects the accuracy of distance-based scoring, reversing the scores of ID and OoD samples, and offer a strategy to overcome this issue. To enhance the explainability and trustworthiness of our method, we also propose a novel attribution technique that explores the latent space to generate attribution maps, highlighting features that deviate from those seen in In-Distribution (ID) classes. To assess the competitiveness of our FFA-based approach, we trained networks on multiple well-known datasets (e.g., KMNIST, EMNIST, CIFAR-10) and evaluated FF-OoD using OoD samples from conventional baselines (e.g., Omniglot, NotMNIST). Additionally, we conducted a qualitative analysis of our interpretability method by manually introducing artifacts into training samples and evaluating the quality of the resulting attribution maps. Our results demonstrate that the latent space of networks trained with FFA naturally separates OoD instances from ID class manifolds, enabling effective OoD detection in spiking neural networks. Furthermore, we showcase the effectiveness of our attribution method in accurately highlighting anomalies injected into input images, providing an intuitive and human-interpretable way to identify the regions responsible for pushing a sample outside its class distribution.
Limitations and Future Work
To obtain a faithful representation of the latent space, our method requires a large number of latent representatives from each class, making it memory-intensive when dealing with datasets with many classes. Our primary focus is on developing mechanisms to reduce intra-class distance, thereby creating a more tractable geometry, especially for instances with small ID datasets. Another key direction for future work stems from the selectivity observed in these networks, which could be leveraged for lifelong learning tasks. We aim to develop a detection mechanism to identify when novel clusters of OoD data emerge and to label them as novel classes.
Acknowledgements
The authors thank the Basque Government for its funding support via the consolidated research groups MATHMODE (ref. T1256-22) and D4K (ref. IT1528-22), and the BEREZ-IA ELKARTEK project (ref. KK-2023/00012). Erik B. Terres-Escudero acknowledges the funding support from TECNALIA Research & Innovation, provided to employees pursuing their doctoral thesis.
Appendix A Reversal of Order in FF-OoD